OpenStudy (anonymous):

HELP MEDAL FOR ANSWER!!!!!!!!!!!!!!!!!!!!!! Which inequality models this problem? The length of a rectangle is four times its width. If the perimeter is at most 130 centimeters, what is the greatest possible value for the width? A. 2w + 2 • (4w) < 130 B. 2w + 2 • (4w) > 130 C. 2w + 2 • (4w) ≤ 130 D. 2w + 2 • (4w) ≥ 130

OpenStudy (anonymous):

PLEASE HELP!

OpenStudy (anonymous):

Well we know that the perimeter is two lengths and two widths: 

OpenStudy (anonymous):

If the maximum is 130, then that means: \[2(length) + 2(width)\le130\]

OpenStudy (anonymous):

Can you go from here?

OpenStudy (anonymous):

NOT REALLY IM BAD AT MATH

OpenStudy (anonymous):

SOMEONE ELSE PLEASE HELP

OpenStudy (anonymous):

length = 4w width = w Plug those into the above equation..

OpenStudy (jdoe0001):

\(\bf length =l\qquad width=w \\ \quad \\ \textit{The length of a rectangle}\implies l\\ \textit{is four times its width.}\implies =4\times w\implies {\color{blue}{ 4w}} \\ \quad \\ \textit{perimeter of a rectangle }=l+w+l+w\implies 2l+2w \\ \quad \\ \textit{perimeter is at most 130 centimeters}\implies perimeter = p \le 130\ cm\\ ------------------------------\\ p=2l+2w\implies p \le 130\ cm\implies 2l+2w \le 130 \\ \quad \\ 2({\color{blue}{ 4w}})+2w \le 130\)

OpenStudy (jdoe0001):

something is AT MOST meaning, it can be THAT MUCH or less, but not above that, thus LESS THAN OR EQUALS TO or \(\large \le\)